OpenStudy (anonymous):

What is the value of a? 6(a+2)+3=3(10-3a)

OpenStudy (openstudier):

First, simplify both sides:\[6a + 12 + 3 = 30 - 9a\]Then, more simplifying:\[6a + 15 = 30 - 9a\]Do you think you can take it from there?

OpenStudy (anonymous):

Not really I really do not understand this

OpenStudy (openstudier):

Ok, so in that first step, I used the distributive property. Do you know what that is?

OpenStudy (anonymous):

Yes

OpenStudy (openstudier):

So, that is how I got that first part. I just distributed on both sides.

OpenStudy (anonymous):

So I do 15 And 30 The Same Way?

OpenStudy (openstudier):

Not exactly. I left it to where you have to move your "a-terms" to one side and move your constants (regular numbers) to the other side.